OpenStudy (adamaero):

Am I starting right: Show lim...(1+x/n)^n = e^x

OpenStudy (adamaero):

Question: Show that the \[\lim_{n \rightarrow \infty}(1 + \frac{ x }{ n })^n = e^x (x>0)\]

OpenStudy (zarkon):

x can actually be any number

OpenStudy (adamaero):

lny = lim... ln(1+(x/n)^n y'/y = n(lim... ln1 + (ln(x) - ln(n)) = n(lim... (1/x) + (1/x) - (n'/n))

OpenStudy (adamaero):

lim... means \[\lim_{n \rightarrow \infty}\]

OpenStudy (zarkon):

you might want to try that again

OpenStudy (adamaero):

am I just suppose to find a definition for e where the whole definition is to the x (or what)?

OpenStudy (adamaero):

so there is no logarithmic differentiation?

OpenStudy (zarkon):

there is...you did it wrong

OpenStudy (adamaero):

OK, I'll try again later. Thanks

OpenStudy (zarkon):

I assume you are trying to use L'Hospitals rule (you should be). Remember that only works with 0/0 or infinity /infinity